4. Define a function f:N → Z by tof n/2 if n is even 1-(n + 1)/2 if n is odd. f(n) = Show that f is a bijection.11 ] 7. Let X = R XR and let R be a relation on X defined as follows ((x,y),(w,z)) ER 4 IC ER\ {0} (w = cx and z = cy.) Is R
Answer #1

Ans. 4. fin - 2 f(n) = 2 - (ne) if n is even if n is oda To show f is bijection, first we show it one-one let f(n) = f (m2) m- 29 sol (n) = 2m il n is add y = -(441) a n--ay-1 = f(n) = -2n-1 - -an- به لم -mr , if n is old f is defined for every n eAus. 7 et x= R XR and R be a relation on x. ((a, g, (w, 2)) ER 7 CEFR 1203 soto (is for Reflexive (cang), (49) nax, gay- (ny)claim :- (Conius, (o, ad) ER. . : (Lug), (0, 2)) ((w, 2), (1, 2)) ER 3 wca zocy and - 6,w, quar - Pa (en) , qa alcy po cica ,Date: Notes Ame. 8. let R be an eavivalance relation on X. then {ado = [6] claim of a, b et it be [ada let & & [a]e and given